Exercise 13.76: Problems - Molarity Part A A laboratory procedure calls for making 550.0 mL of a 1.5 M KNO3 solution How much KNOs in grams is needed? Express your answer using two significant figures Submit My Answers Give Up
Answer #1

volume , V = 550.0 mL

= 0.55 L

we have below equation to be used:

number of mol,

n = Molarity * Volume

= 1.5*0.55

= 0.825 mol

Molar mass of KNO3 = 1*MM(K) + 1*MM(N) + 3*MM(O)

= 1*39.1 + 1*14.01 + 3*16.0

= 101.11 g/mol

we have below equation to be used:

mass of KNO3,

m = number of mol * molar mass

= 0.825 mol * 101.11 g/mol

= 83.42 g

Answer: 83 g

Feel free to comment below if you have any doubts or if this answer do not work. I will edit it if you let me know